OpenStudy (purple_pink):

A grocery store is selling 5 lb of oranges for $4.45. How much will 12 lb of oranges cost? Write a porportion.

OpenStudy (purple_pink):

@KyanTheDoodle

OpenStudy (kyanthedoodle):

What is 12 divided by 5?

OpenStudy (purple_pink):

12/5=2.4

OpenStudy (kyanthedoodle):

OK now just multiply 4.45 by 2.4

OpenStudy (purple_pink):

10.68

OpenStudy (kyanthedoodle):

So that's how much it would be for 12 pounds. The proportion would be 4.45/5 = 10.68/12

OpenStudy (purple_pink):

ohhh i'm kinda getting it

OpenStudy (purple_pink):

1 more please

OpenStudy (kyanthedoodle):

OK